40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Dec 17 00:30:05 2010, Koji, Update, IOO, ITMY / SRM / BS / PRM OPLEVs aligned 
    Reply  Fri Dec 17 03:37:47 2010, rana, Update, SUS, ITMY / SRM / BS / PRM OPLEVs aligned 
Message ID: 4066     Entry time: Fri Dec 17 00:30:05 2010     Reply to this: 4069
Author: Koji 
Type: Update 
Category: IOO 
Subject: ITMY / SRM / BS / PRM OPLEVs aligned 

[Steve and Koji]

The invac OPLEV mirrros were aligned before we get to the PMA party.

The OPLEV mirrors were adjusted in accordance with the optical layout.
Surprisingly the optical layout was enough precise such that we have the healthy red beams on the optical tables.
Steve placed the apertures at the position of the returning spots while I shook the stack to check if the range of the spot motion is sufficient.

The sole thing that has been deviated from the optical layout was that the SRM returning beam had to be reroute
as the SRM has better reflectivity on the AR surface in stead of the HR one.

ELOG V3.1.3-